OpenStudy (anonymous):

how would you get the derivative of xy+2x+3x^2=1 using implicit differentiation

OpenStudy (anonymous):

you use the chaine rule on every y you see... or more easily add y' to your y derivative

OpenStudy (anonymous):

Use y = 1/x - 2 - 3x

OpenStudy (anonymous):

so...your answer should look like 1(y')+2+6x=1...y'=-1-6x

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

do product rule for xy

OpenStudy (anonymous):

= 0, not 1

OpenStudy (anonymous):

y' + 2 + 6x = 0, so y' = -6x - 2

OpenStudy (anonymous):

xy'+y+2+6x=0....xy'=-6x-2....y'=(-6x-2)/x, if I'm not mistaken

OpenStudy (anonymous):

My bad - My answer is wrong

OpenStudy (anonymous):

MathTy is correct. The chain differential of xy is xy' + y

OpenStudy (anonymous):

teamwork, anyways that should be your answer, doing math on the computer without writting it down can be hard

OpenStudy (anonymous):

ok thank you everyone for your help